Date: Thu, 12 Nov 1998 11:12:16 -0800 (PST)
To: 2mass
Subject: IPAC 2MASS WG Meeting #170 Minutes
Cc: chas, stiening, bgreen
Mime-Version: 1.0
Content-Transfer-Encoding: 7bit
Content-MD5: aOESDH0UM4twsuAo8+WPmw==

           IPAC 2MASS Working Group Meeting #170 Minutes

Attendees: R. Cutri, S. Van Dyk, R. Beck, T. Evans, R. Hurt, G. Kopan,
           H. McCallon, R. Tam, W. Wheaton, B. Nelson, L. Fullmer,
           T. Jarrett, D. Kirkpatrick, J. Fowler, D. Engler, J. White


1.) Project Update 

2.) Sampler and Spring release

3.) New and Standing Action Items 


1.) Project Update 

     B. Nelson, fresh from his nuptials, will be advancing on Nov 9 from lowly
postdoc to IPAC sci staff, joining the cut QA team.  Also, on Nov 9, Jeonghee
Rho will join the 2MASS team as a data products analyzer.

     The Leach electronics have been installed and tested in the north.  They
appear to result in data that are "pretty much indistinguishable" from the data
with the old Gatir electronics, however, some subtleties may arise.  But there
was no alternative to making the switch.

     The QA report for the last of the data from 1997 has been written, just in
time for dino to come online next week and increase the production rate by 50%.

     The first night of the Leach electronics (981023n) should define the 
start of a new hardware break-point for the northern operations.  The
break-point should apply to dark and responsivity frames, but not PSFs.

     Pressure is also on the database loads, especially getting loads in without
trouble.  T. Evans reports that a new method she is using appears to be
decreasing the error rate.  DBMAN cannot run at the same time to load north and
south data.

     The 2MASS ERB meeting is coming up this Friday in Amherst.  Attending from
IPAC will be R. Cutri, T. Chester, D. Kirkpatrick, B. Green, and T. Handley.  
R. Cutri has circulated a URL with the meeting agenda.  Three questions will be
posed and hopefully answered in full:  1) Is the project on track to meet the 
Level 1 specs; 2) is the project still timely; and, 3) are the initial data 
release plans realistic and appropriate?

     Anomalies of the week are highlighted again by the J-banding in the south.
D. Kirkpatrick has been spot-checking scans.  He finds that the problem is seen
in all the sci scans he has looked at, at a level of a few-tenths of a DN, to
as bad as a full DN.  G. Kopan added that it is mostly seen on the west side of
the scans; usually the east edge was not affected.  The flattened frames showed 
the northwest quad to be the most affected, with a probable sinusoid impressed 
on the background.  G. Kopan finds that the J band images from 980319s onward 
started showing periodic bands (90--110") aligned with cross-scan that varied 
in amplitude and somewhat in spatial period.  Since the spatial period is nearly
the same as the in-scan step size, the wave adds constructively in the coadds. 
The amplitude varies on timescales often faster than the background offset
stack size (42 frames), so DFLAT only partially removes the effect. 

     The possibility of generating an objective measure of the effect for QA
is being investigated. It is unknown at this point how much of the data from the
south are affected.  R. Beck has run some sci scans in recent southern nights 
for analysis.  For additional information, see  The problem
is definitely bad for galaxy photometry in GALWORKS, according to T. Jarrett.

     Joining the list of the week's anomalies are hot pixels seen in the south,
which are not being handled by FREXAS.  They are not easy to identify and to
mask.  R. Cutri, by eye, states that the pixel seems to be the same one in all
data.  The hot pixels are seen in H and K.

     Finally, as H. McCallon reported, there is the distortion in the southern
cameras.  K has a lot more distortion than J and H.  A few nodes can even be
seen in the K "pincushion".  This is clearly a problem that needs to be 
corrected in 2MAPPS v3.0.

     R. Cutri asked whether the new image moment ratio diagnostic had made its
way yet into QA.  R. Hurt reported that it had, as of the 24 QA on Monday, 
although the QA webpage was still providing the old aspect ratio vs. positional
angle plot.  R. Hurt needed to get the necessary cgi script running to generate
the new plot.  He will consult with L. Fullmer to bring this about.

     R. Cutri asked of L. Fullmer why the scan overlap plots were not in the
QA nightly page.  She reported that she had fixed it, so that the plots should 
now be appearing.  (She was trying to make a graphical change to distinguish
photometric from nonphotometric points in the plots.)

     R. Tam is checking the bright star artifacts in the Sampler night.  He has
flagged all the diffraction spike artifacts.  These can now be cleaned out of
the Sampler.  He reported on wide diagonal bands that can be seen in some J 
scans.  But, it was agreed that these scans are near the scan with Beta Pegasi
(Scheat), and are therefore reflection artifacts from that bright star.  As G.
Kopan and S. Van Dyk offered, similar artifacts are seen in the IC 443 mosaic
resulting from the presence of eta Gemini in a nearby scan.

     T. Jarrett is working on a journal paper describing the extended source
catalog.  He hopes to have it done (in press?) by the spring data release.     

2.) Sampler and Spring release

     The Sampler release date is no longer Nov 3, but shortly thereafter.  
Hopefully, no more than a week later....  [which was wishful thinking]

     The spring release will consist of data from 970606n to 971203n, plus
980106n to 980129n.  It will include about 2900--3500 square degrees of sky,
depending on the quality allowed.  It was originally to cover the +12 to +18
dec band, but will probably not be strictly this, since this band has not been
filled with quality=10 data.  So, the release area will likely stray from this.

     40% of data are now quality 3 or below.  Quality 0 data are immediately
reobserved, while quality 1--2 data are not as readily reobserved.  Quality 3
and above data will go into the catalogs.  The numbers of photometric nights, so
far, are pretty much at historical values.  At the current rate, the survey 
should be finished at the end of 2000; more likely a little beyond this point.

3.) New and Standing Action Items

%                                                      %
%                 ACTION ITEMS                         %
%                                                      %

      R. Cutri did not assign any new action items.  

(Standing Action Items:)

a) Team members should continue testing the IRSA tools.

b) Team members should still contemplate the sci-staff structure proposals and 
voice their opinions.

c) Team members should continue to analyze data for spring data release.

d) Team members are reminded that performance reports covering October and 
November will be due on Dec 2.